OpenStudy (anonymous):

What is the length of the base of a triangle with height 20 cm and area 150 cm2? (Points : 2) 3.75 cm 10 cm 15 cm 7.5 cm @jim_thompson5910

jimthompson5910 (jim_thompson5910):

A = (b*h)/2 150 = (b*20)/2 solve for b

jimthompson5910 (jim_thompson5910):

if you're stuck, multiply both sides by 2 and flip both sides to get 20b = 300

OpenStudy (anonymous):

15 cm

jimthompson5910 (jim_thompson5910):

yep